# Informed Asymmetric Actor-Critic: Leveraging Privileged Signals Beyond Full-State Access

### v1

@@ 4.1 Informed History-based Value Functions
Hence, by Lemma 4.3, the informed history value function $V^{\pi}(h,i)$ is an unbiased estimator of the standard history value $V^{\pi}(h)$. Put differently, $V^{\pi}(h,i)$ provides, in expectation, the same signal as the standard history value function.
In the special case where the privileged information $i$ corresponds to the full state $s\in\mathcal{S}$, i.e., $i=s$, the informed history value function reduces to the history-state value function of Baisero and Amato [10] (cf. Corollary A.1).
@@ 4.2 Informed Asymmetric Policy Gradient
Based on the informed history-based Q-function, we define the informed asymmetric policy gradient as:
$\nabla_{\theta}^{\text{IAAC}}J(\pi_{\theta})=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}Q^{\pi}(h_{t},i_{t},a_{t})\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid h_{t})\right],$
where the actor conditions on observable histories, and the critic may condition on additional privileged inputs. We show that this gradient remains unbiased:
@@
Given Equation 1 and following the Lemmas 4.2-4.3, we have
[…]
In (a) and (d), we use the linearity of the expectation to decompose or combine the summation over $t$ and the expectation over $(h_{t},i_{t},a_{t})$, respectively. In (b), using Lemma 4.2, we substitute $Q^{\pi}(h,a)$ with $\mathbb{E}_{i\mid h}\left[Q^{\pi}(h,i,a)\right]$, as the informed history-action value function is an unbiased estimate of $Q^{\pi}(h,a)$. By applying the law of total expectation in (c), i.e., $\mathbb{E}_{h_{t},a_{t},i_{t}}\left[\cdot\right]=\mathbb{E}_{h_{t},a_{t}}\left[\mathbb{E}_{i_{t}\mid h_{t}}[\cdot]\right]$, we can rewrite the expression. This concludes the proof. ∎
Thus, the critic can incorporate additional training-time signals without biasing policy updates. This result generalizes prior work and recovers the asymmetric policy gradient presented by Baisero and Amato [10] for $i_{t}=s_{t}$ (cf. Corollary A.2).
Based on Theorem 4.1, we introduce the informed history critic $\hat{V}$$\mathcal{H}\times\mathcal{I}\to\mathbb{R}$, which estimates the informed history value $V^{\pi}(h_{t},i_{t})$ given the history $h_{t}$ and additional information $i_{t}$. Combined with a history-dependent policy model $\pi_{\theta}(a_{t}\mid h_{t})$, this yields an asymmetric actor-critic method, which we refer to as informed asymmetric actor-critic (IAAC). The informed asymmetric policy gradient is approximated by sampling $\hat{\nabla}_{\theta}^{\text{IAAC}}J(\pi_{\theta})~=~\mathbb{E}\left[\sum_{t}\gamma^{t}\ \delta_{t}\ \nabla_{\theta}\log\pi_{\theta}(a_{t}\mid h_{t})\right],$ where the TD errors $\delta_{t}=r_{t}+\gamma\ \hat{V}(h_{t+1},i_{t+1};\vartheta)-\hat{V}(h_{t},i_{t};\vartheta)$ are computed using the critic’s informed value estimates.
@@ 5 INFORMATIVENESS OF PRIVILEGED SIGNALS
While the informed asymmetric actor-critic framework guarantees unbiased policy gradients for arbitrary additional information (cf. Theorem 4.1), not all privileged signals are equally beneficial for learning. In practice, some signals may accelerate policy optimization, while others may degrade value estimation or introduce instability when poorly correlated with the environment’s true state.
This motivates the need to quantify the informativeness of additional information $i_{t}\in\mathcal{I}$ with respect to the underlying control task. Specifically, we seek criteria that assess whether $i_{t}$ provides useful structure for value estimation and can improve the policy learning of actor-critic methods.
In the following, we formalize two criteria that enable the comparison of different forms of privileged signals in terms of their informativeness about true returns.
5.1 Informativeness via Kernel-based Independence Criterion
The relevance of a privileged signal in the critic depends on whether it captures information about the return that is not already encoded in the history of past observations and actions. To quantify this, we evaluate the conditional dependence between the return $G_{t}:=\sum_{j=0}^{T-t-1}\gamma^{j}R_{t+j}$and , conditioned on $h_{t}$ and $a_{t}$, using a non-parametric kernel-based approach.
Kernel methods embed probability distributions into a reproducing kernel Hilbert space (RKHS), where statistical relationships can be expressed as distances between mean elements. For istance, a distribution $\mathbb{P}$ over $X$ is represented as the kernel mean embedding $\mu_{\mathbb{P}}:=\mathbb{E}_{x\sim\mathbb{P}}[k(x,\cdot)]\in\mathscr{H}_{X}$, where $k(\cdot)$ denotes a positive-definite kernel function.
Kernel-based measures such as Maximum Mean Discrepancy (MMD) [24] and the Hilbert-Schmidt Independence Criterion (HSIC) [25] enable non-parametric tests for differences in distributions and (in)dependence, respectively. In this work, we adopt the Hilbert-Schmidt Conditional Independence Criterion (HSCIC) [26], which extends HSIC to the conditional setting.
Given random variables $X,Y,Z$, HSCIC measures the RKHS norm between the conditional joint embedding and the product of conditional marginals:
$\mathscr{I}(X,Y\mid Z):=\left\|\mu_{\mathbb{P}_{X,Y|Z}}-\mu_{\mathbb{P}_{X|Z}}\otimes\mu_{\mathbb{P}_{Y|Z}}\right\|_{\mathscr{H}_{X}\otimes\mathscr{H}_{Y}}.$
Under standard conditions, such as a characteristic kernel on $X\times Y$ and that the conditional probability distributions given $Z$ admit a regular version, HSCIC equals zero almost surely if and only if $X\perp\!\!\!\perp Y\mid Z$ (cf. Theorem C.1).
We leverage HSCIC to test whether $i_{t}$ is conditionally dependent on the return $G_{t}$ given $h_{t}$ and $a_{t}$. This analysis does not require a trained critic and can be applied to trajectories collected under a random or exploratory policy. As a result, informativeness can be estimated prior to training, enabling informed decisions regarding which auxiliary signals to include in the critic.
Let $X_{t}=G_{t}$, $Y_{t}=i_{t}$, and $Z_{t}=(h_{t},a_{t})$. Given samples $\{(x_{i},y_{i},z_{i})\}_{i=1}^{n}$, the empirical plug-in estimator of the squared HSCIC is given by
$\begin{split}\hat{\mathscr{I}}^{2}_{X,Y\mid Z}(\cdot)=k^{\top}_{Z}W(K_{X}\odot K_{Y})W^{\top}k_{Z}&-2k^{\top}_{Z}W((K_{X}W^{\top}k_{Z})\odot(K_{Y}W^{\top}k_{Z}))\\ &+(k^{\top}_{Z}WK_{X}W^{\top}k_{Z})(k^{\top}_{Z}WK_{Y}W^{\top}k_{Z}),\end{split}$
where $k_{Z}$ denotes the kernel function $k_{Z}(\cdot)$ on $Z$ evaluated at the estimator input $z_{i}$$K_{Y}]_{ij}=k_{Y}(y_{i},y_{j})$, $W=(K_{Z}+\lambda\mathbf{I})^{-1}$, and $\odot$ denotes the element-wise product.
Hypothesis test.
We can construct a hypothesis test based on permutation testing to assess statistical significance. Let
$\mathbb{H}_{0}:G_{t}\perp\!\!\!\perp i_{t}\mid h_{t},a_{t},\quad\text{and}\quad\mathbb{H}_{1}:G_{t}\not\!\perp\!\!\!\perp i_{t}\mid h_{t},a_{t}.$
To obtain a scalar test statistic, we average the pointwise HSCIC values over all sample realizations $z_{i}$ of $Z$:
$\bar{\mathscr{I}}_{X,Y\mid Z}=\frac{1}{n}\sum_{i=1}^{n}\hat{\mathscr{I}}_{X,Y\mid Z}(z_{i}).$
Under $\mathbb{H}_{0}$, the null distribution is generated by independently permuting $\{y_{i}\}$ while keeping $\{x_{i},z_{i}\}$ fixed. For each of $B$ permutations, we compute $\bar{\mathscr{I}}^{(b)}$, and evaluate the empirical $p$-value as $p=\frac{1}{B}\sum_{b=1}^{B}\mathbb{I}\left[\bar{\mathscr{I}}^{(b)}\geq\bar{\mathscr{I}}\right]$, where $\mathbb{I}$ denotes the indicator function. We reject $\mathbb{H}_{0}$ at significance level $\alpha\geq 0$ if $p<\alpha$, indicating that $i_{t}$ is informative for predicting $G_{t}$ given $h_{t}$ and $a_{t}$.
5.2 Informativeness via Return Prediction Error
Beyond pre-training evaluation of privileged signals using HSCIC, we propose a complementary post-hoc criterion to quantify the utility of a privileged signal $i_{t}$ in improving return prediction after training. Importantly, this evaluation focuses exclusively on the value prediction accuracy and does not involve any policy-related metrics.
Consider two critics trained on the same task: a symmetric critic $Q(h_{t},a_{t})~=~\mathbb{E}[G_{t}\mid h_{t},a_{t}]$, and an asymmetric critic $Q(h_{t},i_{t},a_{t})~=~\mathbb{E}[G_{t}\mid h_{t},i_{t},a_{t}]$.
Let $T$ be the length of a trajectory. We define the pointwise reduction in squared error from conditioning on $i_{t}$ as
$E_{t}:=\left(Q(h_{t},a_{t})-G_{t}\right)^{2}-\left(Q(h_{t},i_{t},a_{t})-G_{t}\right)^{2},$ (8)
where a positive $E_{t}$ indicates improved return prediction at time $t$. To summarize the benefit over a trajectory, we compute the empirical mean and variance of $\{E_{t}\}_{t=0}^{T}$:
$\hat{x}=\frac{1}{T}\sum_{t=0}^{T-1}E_{t},\quad\hat{\sigma}^{2}=\frac{1}{T}\sum_{t=0}^{T-1}\left(E_{t}-\hat{x}\right)^{2}.$
Definition 5.1 ($(\epsilon,\delta)$-Informativeness).
A privileged signal $i_{t}$ is said to be $(\epsilon,\delta)$-informative if, with probability at least $1-\delta$, the expected gain satisfies
$\mathbb{E}[E_{t}]\geq\epsilon:=\hat{x}-\sqrt{\frac{2\hat{\sigma}^{2}\log(2/\delta)}{T}}-\frac{2C\log(2/\delta)}{3T},$ (9)
where $C:=\max_{0\leq t<T}|E_{t}|$.
This bound, derived via Bernstein’s inequality, provides a lower confidence bound on the expected gain from including $i_{t}$, serving as both a quantitative metric and a test statistic for a hypothesis test.
Hypothesis test.
We test the hypotheses
$\mathbb{H}_{0}:\epsilon\leq 0,\quad\text{and}\quad\mathbb{H}_{1}:\epsilon>0,$
where $\epsilon$ is the lower bound defined in Defintion 5.1. We reject $\mathbb{H}_{0}$ if $\epsilon>0$, concluding that the privileged signal $i_{t}$ is $(\epsilon,\delta)$-informative with confidence $1-\delta$. Otherwise, we fail to reject $\mathbb{H}_{0}$, indicating insufficient evidence that $i_{t}$ improves value prediction. Hence, this test offers a statistically grounded post-training criterion for evaluating privileged signals based on their empirical effect on return prediction accuracy.

### v2

@@ 3.1 Informed History-based Value Functions
As discussed by Baisero and Amato [4], the history $h_{t}$ determines the agent’s future behavior under a fixed policy, while the state governs the environment’s rewards and dynamics. In our setting, the privileged signal $i_{t}$ provides additional state-dependent context that can reduce the ambiguity about the underlying state given a fixed history. Formally, letting $H(s_{t}|h_{t})$ denote the conditional entropy of the state, standard information-theoretic results imply $E_{i_{t}|h_{t}}~\left[H(s_{t}|h_{t},i_{t})\right]~\leq~H(s_{t}|h_{t})$, so conditioning on $i_{t}$ never increases, and on average reduces, uncertainty about the true state. This does not guarantee that $V^{\pi}(h_{t},i_{t})$ is inherently easier to approximate than $V^{\pi}(h_{t})$; rather, it is the conditional expectation of a return random variable whose variance may be lower. In particular, by the law of total variance applied to the return $G_{t}$, $E_{i_{t}|h_{t}}~\left[\text{Var}(s_{t}|h_{t},i_{t})\right]~\leq~\text{Var}(s_{t}|h_{t})$, so appropriate choice of $i_{t}$ can reduce the variance of value targets.
When the privileged signal coincides with the full state, $i_{t}=s_{t}$, our formulation reduces to the history-state value function of Baisero and Amato [4] (cf. Corollary A.1 in Appendix A), showing that their framework is a special case of our more general formalization.
@@ 3.2 Informed Asymmetric Policy Gradient
Using the informed history-based $Q$-function, we define the informed asymmetric policy gradient
$\nabla_{\theta}^{\text{IAAC}}J(\pi_{\theta})=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}Q^{\pi}(h_{t},i_{t},a_{t})\nabla_{\theta}\log\pi_{\theta}(a_{t}|h_{t})\right],$
where the policy $\pi_{\theta}$ depends only on the history $h_{t}$, while the critic may additionally condition on training-time information $i_{t}$.
Incorporating privileged state-conditioned information into the critic does not bias the policy gradient: for any choice of $i_{t}$, the informed asymmetric policy gradient coincides with the standard policy gradient.
@@
Given Equation 1 and Lemma 3.2, we have
$\nabla_{\theta}J(\pi_{\theta})$ $=\mathbb{E}\Bigg[\sum_{t}\gamma^{t}\ Q^{\pi}(h_{t},a_{t})\ \nabla_{\theta}\log\pi_{\theta}(a_{t}|h_{t})\Bigg]$
$\overset{\text{(a)}}{=}\sum_{t}\gamma^{t}\ \mathbb{E}_{h_{t},a_{t}}\Big[Q^{\pi}(h_{t},a_{t})\ \nabla_{\theta}\log\pi_{\theta}(a_{t}|h_{t})\Big]$
$\overset{\text{(b)}}{=}\sum_{t}\gamma^{t}\ \mathbb{E}_{h_{t},a_{t}}\Big[\mathbb{E}_{i_{t}}\big[Q^{\pi}(h_{t},i_{t},a_{t})\big|h_{t}\big]\ \nabla_{\theta}\log\pi_{\theta}(a_{t}|h_{t})\Big]$
$\overset{\text{(c)}}{=}\sum_{t}\gamma^{t}\ \mathbb{E}_{h_{t},i_{t},a_{t}}\Big[Q^{\pi}(h_{t},i_{t},a_{t})\ \nabla_{\theta}\log\pi_{\theta}(a_{t}|h_{t})\Big]$
$\overset{\text{(d)}}{=}\mathbb{E}\Bigg[\sum_{t}\gamma^{t}\ Q^{\pi}(h_{t},i_{t},a_{t})\ \nabla_{\theta}\log\pi_{\theta}(a_{t}|h_{t})\Bigg]$
$=\nabla_{\theta}^{\text{IAAC}}J(\pi_{\theta}).$
Here, (a) and (d) follow from linearity of expectation, (b) uses Lemma 3.2, and (c) applies the law of total expectation. This concludes the proof. ∎
This result strictly generalizes prior work on asymmetric actor-critic methods. In particular, it recovers the asymmetric policy gradient of Baisero and Amato [4] as a special case when $i_{t}=s_{t}$ (cf. Corollary A.2). Hence, Theorem 3.4 extends asymmetric actor-critic theory beyond state-based critics and provides a formal justification for the use of arbitrary state-conditioned training-time signals.
Informed history critic.
As in symmetric actor-critic methods, we learn a value function using TD learning, and we use the TD error to form a low-variance advantage estimate. The informed history critic $\hat{V}$$\mathcal{H}~\times~\mathcal{I}~\to~\mathbb{R}$ additionally receives a privileged signal and approximates the informed history value $V^{\pi}(h_{t},i_{t})$. Combined with a history-dependent actor $\hat{\pi}_{\theta}(a_{t}|h_{t})$, this yields the informed asymmetric actor-critic (IAAC). The resulting policy gradient estimator is
$\hat{\nabla}_{\theta}^{\text{IAAC}}J(\pi_{\theta})~=~\mathbb{E}\left[\sum_{t}\gamma^{t}\ \hat{A}^{\text{IAAC}}_{t}\ \nabla_{\theta}\log\hat{\pi}_{\theta}(a_{t}|h_{t})\right],$
where $\hat{A}^{\text{IAAC}}_{t}$ is computed using the informed history critic.
@@ 4 Informativeness of Privileged Signals
While the informed asymmetric actor-critic framework guarantees unbiased policy gradients for any state-conditioned privileged signals (cf. Theorem 3.4), not all such signals are equally useful for learning in practice. The choice of $i_{t}$ can affect the critic’s learning efficiency and stability, motivating the need for criteria to select the most informative signals.
Here, we focus on the informativeness of $i_{t}\in\mathcal{I}$ from the perspective of predicting future returns, which is the critic’s primary role. To this end, we introduce two criteria that quantify the informativeness of privileged signals and enable systematic comparison across different choices of $i_{t}$.
4.1 Residual Informativeness
The utility of a privileged signal $i_{t}$ in an informed critic depends on whether it provides information about future returns beyond what is already encoded in the observable history-action pair $(h_{t},a_{t})$. At the population level, this corresponds to the conditional independence (CI) hypothesis
$\mathbb{H}_{0}^{\text{CI}}:G_{t}\perp\!\!\!\perp i_{t}|h_{t},a_{t},$
which factorizes the joint distribution as
$p(G_{t},i_{t},h_{t},a_{t})=p(h_{t},a_{t})p(G_{t}|h_{t},a_{t})p(i_{t}|h_{t},a_{t}).$
Rejecting $\mathbb{H}_{0}^{\text{CI}}$ indicates that $i_{t}$ provides additional, non-redundant information about future returns.
One typically requires samples from both $\mathbb{H}_{0}^{\text{CI}}$ and $\mathbb{H}_{1}^{\text{CI}}$ to perform this test. Unfortunately, directly obtaining samples following $\mathbb{H}_{0}^{\text{CI}}$ is generally infeasible. Instead, interaction with the environment yields samples from the unknown joint distribution $p(G_{t},i_{t},h_{t},a_{t})=p(h_{t},a_{t})p(G_{t},i_{t}|h_{t},a_{t})$, where CI cannot be assumed. Sampling independently from $p(G_{t}|h_{t},a_{t})$ would require learning this conditional distribution, which is challenging for high-dimensional histories, and risks introducing noise into any direct test.
To circumvent this, we first remove the predictable components of $G_{t}$ and $i_{t}$ explained by $(h_{t},a_{t})$ and test for dependence in the residuals. As we explain below, testing residual independence provides insight into the independence of the original random variables, while simultaneously allowing sampling from a tractable approximation of $p(G_{t}|h_{t},a_{t})$.
Definition 4.1 ($\alpha$-residual informativeness).
Let $G_{t}$, $i_{t}$, $h_{t}$, and $a_{t}$ be random variables with finite second moments. Define the conditional-mean residuals
$e_{G_{t}}:=G_{t}-\mathbb{E}[G_{t}|h_{t},a_{t}],\qquad e_{i_{t}}:=i_{t}-\mathbb{E}[i_{t}|h_{t},a_{t}].$
A privileged signal $i_{t}$ is $\alpha$-residual-informative if a statistical test of independence rejects the null hypothesis
$\mathbb{H}_{0}^{\text{res}}:e_{G_{t}}\perp\!\!\!\perp e_{i_{t}},$
in favor of the alternative $\mathbb{H}^{\text{res}}_{1}:e_{G_{t}}\not\!\perp\!\!\!\perp e_{i_{t}}$, at significance level $\alpha>0$.
CI implies residual independence, $\mathbb{H}_{0}^{\text{CI}}\Rightarrow\mathbb{H}_{0}^{\text{res}}$, but the converse need not hold in general [45]. Hence, the residual test provides a necessary, but not sufficient, condition for CI, which aligns with our goal of detecting whether $i_{t}$ carries additional predictive information about $G_{t}$ beyond $(h_{t},a_{t})$. This reduces the problem to testing whether the unexplained components of $G_{t}$ and $i_{t}$ remain statistically dependent.
Testing $\mathbb{H}_{0}^{\text{res}}$ ideally requires samples from
$p(e_{G_{t}}|h_{t},a_{t})\,p(e_{i_{t}}|h_{t},a_{t})\,p(h_{t},a_{t}),$
which, as with the original random variables, is generally intractable. To approximate a sample of $G_{t}$ independent of $i_{t}$ but with the correct conditional mean, we construct a surrogate $\tilde{G}_{t}$ in two steps: (i) we draw from the marginal distribution $p(G_{t})$ of $G_{t}$, which is readily available from the collected episodes; (2) we shift these samples to preserve the estimated conditional mean $\mathbb{E}[G_{t}|h_{t},a_{t}]$. Thus,
$\tilde{G}_{t}^{\text{null}}=\tilde{G}_{t}-\mathbb{E}[\tilde{G}_{t}]+\mathbb{E}[G_{t}|h_{t},a_{t}].$
By construction, the distribution of this sample exactly matches the first moment of the target distribution $p(G_{t}|h_{t},a_{t})$, while its higher-order moments follow those of the marginal distribution $p(G_{t})$. The corresponding null residual
$e_{\tilde{G}_{t}}^{\text{null}}:=\tilde{G}^{\text{null}}_{t}-\mathbb{E}[G_{t}|h_{t},a_{t}]=\tilde{G}_{t}-\mathbb{E}[\tilde{G}_{t}]$
is therefore independent of $e_{i_{t}}$, satisfying $\mathbb{H}^{\text{res}}_{0}$.
Practical Implementation.
In practice, we approximate the conditional expectations $\mathbb{E}[G_{t}|h_{t},a_{t}]$ and $\mathbb{E}[i_{t}|h_{t},a_{t}]$ using regression models capable of capturing nonlinear relationships (e.g., random forests or neural networks), since $G_{t}$ and $i_{t}$ typically exhibit nonlinear dependence. Histories are generally encoded using a recurrent network to produce a fixed-length representation $z_{t}=f_{\text{RNN}}(h_{t})$, yielding residual estimates
$\hat{e}_{G_{t}}:=G_{t}-\hat{\mathbb{E}}[G_{t}|z_{t},a_{t}],\quad\hat{e}_{i_{t}}:=i_{t}-\hat{\mathbb{E}}[i_{t}|z_{t},a_{t}],$ (7)
and for the null sample, we have $\hat{e}_{\tilde{G}_{t}}^{\text{null}}~:=~\tilde{G}_{t}~-~\hat{\mathbb{E}}[G_{t}|z_{t},a_{t}]$. We quantify the dependence between $\hat{e}_{G_{t}}$ and $\hat{e}_{i_{t}}$ using a general dependency measure $\rho(\cdot,\cdot)$ (e.g., mutual information or Hilbert-Schmidt Independence Criterion). By design, $\rho(\hat{e}_{G_{t}},\hat{e}_{i_{t}})=0$ under independence and increases with stronger dependence. To account for finite-sample variability and temporal correlations, we approximate the surrogate null by episode-wise permutation of the observed residuals. Residuals $\hat{e}_{G_{t}}$ are shuffled across independently collected episodes while preserving the temporal order within each episode, which maintains the estimated conditional mean structure and respects intra-episode dependencies. Cross-fitting is used to prevent overfitting: the dataset is split into $K$ folds, with regressors trained on K1K-1 folds and evaluated on the held-out fold.
We compute the empirical $p$-value with $B$ permutations as
$p=\frac{1+\sum_{b=1}^{B}\mathbf{1}\{\rho(\hat{e}^{(b)}_{G_{t}},\hat{e}^{(b)}_{i_{t}})\geq\rho(\hat{e}_{G_{t}},\hat{e}_{i_{t}})\}}{B+1},$ (8)
where $\hat{e}^{(b)}_{G_{t}},\hat{e}^{(b)}_{i_{t}}$ denote residuals under the $b$-th permutation and $\mathbf{1}\{\cdot\}$ is the indicator function, which equals $1$ if the condition inside the braces holds and 0 otherwise. The privileged signal $i_{t}$ is declared $\alpha$-residual-informative if $p<\alpha$. Algorithm 1 summarizes the procedure.
Notably, the $\alpha$-residual-informativeness criterion can be evaluated on episodes collected under any exploratory policy, without requiring a trained actor or critic. In particular, a random policy can be used, although coverage of the state-action space may be limited. This allows estimating the informativeness of privileged signals even before training and supports informed decisions about which signals to include. To obtain meaningful residuals, the history encoding function $f_{\text{RNN}}(\cdot)$ should be trained in advance to produce effective representations of past observations and actions.
Input: Independence measure $\rho(\cdot,\cdot)$, number of folds $K$, independently collected episodes $\{h_{t},a_{t},i_{t},G_{t}\}_{t=0}^{T-1}$, number of permutations $B$, significance level $\alpha$.
 Encode histories with recurrent network: $z_{t}=f_{\text{RNN}}(h_{t})$.
 Partition data into $K$ folds for cross-fitting
for each fold $k=1$ to $K$ do
  Train regression models on the K1K-1 other folds.
  Predict $\hat{\mathbb{E}}[G_{t}|z_{t},a_{t}]$ and $\hat{\mathbb{E}}[i_{t}|z_{t},a_{t}]$ on the $k$-th fold.
  Compute residuals $\hat{e}_{G_{t}}$ and $\hat{e}_{i_{t}}$ on the $k$-th fold (Eq. 7).
end for
 Aggregate cross-fitted residuals across held-out folds.
 Compute observed dependence $\rho_{\text{obs}}$ using $\rho(\cdot,\cdot)$.
for $b=1$ to $B$ do
  Permute residuals $\hat{e}_{G_{t}}$ on episode-level.
  Compute $\rho^{(b)}_{\text{perm}}$ for permuted residuals.
end for
 Compute empirical $p$-value according to Eq. 8.
Output: Declare $i_{t}$ $\alpha$-residual-informative if $p<\alpha$.
Algorithm 1 $\alpha$-Residual-Informativeness Test
4.2 Informativeness via Return Prediction Gain
Informativeness can also be assessed via the critic’s predictive accuracy: a privileged signal is useful if its inclusion improves value estimation. Hence, this post-hoc criterion provides a quantitative measure of a signal’s utility for return prediction. It only requires episodes collected under any fixed policy and does not depend on policy performance.
Consider a symmetric critic $\hat{Q}(h_{t},a_{t})$ and an informed asymmetric critic $\hat{Q}(h_{t},i_{t},a_{t})$. For a set of $N$ episodes $\{\tau_{j}\}_{j=1}^{N}$, we define the episode-level squared-error gain
$L^{\tau_{j}}:=\frac{1}{T_{j}}\sum_{t=0}^{T_{j}-1}\Big((\hat{Q}(h_{t},a_{t})-G_{t})^{2}-(\hat{Q}(h_{t},i_{t},a_{t})-G_{t})^{2}\Big),$
where $T_{j}$ is the length of episode $\tau_{j}$ and $G_{t}$ the true future return. $L^{\tau_{j}}>0$ indicates that including $i_{t}$ improves return prediction on that episode.
Definition 4.2 ($(\epsilon,\delta)$-prediction informativeness).
A privileged signal $i_{t}$ is said to be $(\epsilon,\delta)$-prediction-informative with $\epsilon\geq 0$ if a statistical test rejects the null hypothesis
$\mathbb{H}_{0}:\mathbb{E}[L^{\tau}]\leq\epsilon,$
in favor of the alternative $\mathbb{H}_{1}:\mathbb{E}[L^{\tau}]>\epsilon$, at significance level $\delta$, based on the set of episode-level gains $\{L^{\tau_{j}}\}_{j=1}^{N}$.
Practical Implementation.
In practice, the test procedure depends on the number of collected episodes $N$. For small sample sizes, we apply a one-sided bootstrap test: we resample $B>N$ episodes with replacement, compute bootstrap means $\{\bar{L}^{*}_{b}\}_{b=1}^{B}$, and estimate the empirical $p$-value as
$p=\frac{1+\sum_{b=1}^{B}\mathbf{1}\{\bar{L}^{*}_{b}\leq 0\}}{B+1},$
rejecting $\mathbb{H}_{0}$ if $p<\delta$. For larger sample sizes (e.g., N>1, [^0] N>1{,}000), which we apply an one-sided $t$-test based on the sample mean $\bar{L}$ and variance $\hat{\sigma}^{2}$. The test statistic is $t~=~\frac{\bar{L}-\epsilon}{\hat{\sigma}/\sqrt{N}}$, which under $\mathbb{H}_{0}$ follows a Student-$t$ distribution with N1N-1 degrees of freedom. We reject $\mathbb{H}_{0}$ if the corresponding one-sided $p$-value is below $\delta$.
4.3 Selecting Privileged Signal Generators
Typically, one needs to select a signal generator from a candidate set $\mathcal{C}=\{I^{(m)}\}_{m=1}^{M}$, with each providing privileged information $i_{t}^{(m)}~\sim~I^{(m)}(\cdot|s_{t})$. Different generators may encode distinct aspects of the state and are not directly comparable a priori.
Both proposed informativeness criteria support principled selection from $\mathcal{C}$ via hypothesis testing. For each $I^{(m)}$, we evaluate the signal sequence $\{i_{t}^{(m)}\}_{0\leq t<T}$ and deem the generator informative if the corresponding null hypothesis is rejected. In the pre-training setting, we apply the $\alpha$-residual-informativeness test and select generators for which residual independence is rejected at significance level $\alpha$. In the post-hoc setting, we apply the $(\epsilon,\delta)$-prediction-informativeness test and retain generators that yield a statistically significant improvement in value prediction accuracy ($\epsilon=0$). Among informative candidates, one may select the generator with the largest effect size (or smallest $p$-value), yielding the signal that provides the strongest additional information about future returns under the respective criterion.